A boy had twice as many sums wrong as right;if he had 24 sums in all;how many of them were wrong?
Answer by Theo(13342)

r = number he got right.
w = number he got wrong.

r + w = 24

this means the number he got right plus the number he got wrong is equal to the total number of sums that he had to solve.

w = 2r

this means that the number he got wrong is twice as many as the number he got right.

since w = 2r, then replace w with 2r in the equation of r + w = 24.

you will get r + w = 24 becomes r + 2r = 24.

combine like terms to get 3r = 24.

divide both sides of the equation by 3 to get r = 8.

since w = 2r, then w = 16.

he got 8 right and 16 wrong.